OpenStudy (anonymous):

what is the value of the digit 3 in the number 23145? a.)3*1 b.)3*5 c.)3*25 d.)3*125

OpenStudy (iluvsoccer):

23, 145 is the number right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thx

OpenStudy (iluvsoccer):

Sorry, my mistake.

OpenStudy (anonymous):

so that not the answer

OpenStudy (welshfella):

Hold on - the base cannot be 5 because the last digit is 5

OpenStudy (anonymous):

what is the base 8 representation of the number 11101112? a.)231 8 b.)329 8 c.)3478 d.)385 8

OpenStudy (er.mohd.amir):

may be Q is 2314 and base is 5 so ans is 3*5^2=3*25

OpenStudy (welshfella):

could be

OpenStudy (anonymous):

ok wat is the answer to the next question

OpenStudy (er.mohd.amir):

167 on base 8

OpenStudy (anonymous):

i have no idea wat the answer is cuz i dont have a book

OpenStudy (er.mohd.amir):

check the Q or option

OpenStudy (anonymous):

what is the base 10 representation of the number 10468? a.)500 b.)520 c.)550 d.)580

OpenStudy (anonymous):

last question

OpenStudy (er.mohd.amir):

Q is 1046 base 8 then ans 550 base 10

OpenStudy (er.mohd.amir):

use abcd base k then ans in base 10 is d*k^0+c*k^1+b*k^3+a*k^3=sum is ans in base 10

OpenStudy (anonymous):

i passed thx for the help